OpenStudy (anonymous):

what is the slope intercept form of f=4(s-65)+10

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Which variable is for the vertical axis, and which one is for the horizontal axis?

OpenStudy (anonymous):

i dont know actually! its a paragraph you have to read

OpenStudy (anonymous):

what do f and s stand for?

OpenStudy (anonymous):

f=find & s=speed but i have tu put the equation in slope intercept form

OpenStudy (anonymous):

Well, it doesn't matter much. Solving one for slope intercept will be correct, it just depends on how you wanted to graph it. But the one you solve for will determine that.

OpenStudy (anonymous):

Hmm, the teacher may not agree with that, even though you are correct.

OpenStudy (anonymous):

@polpak & @michael.3245 do u guys have an answer

OpenStudy (anonymous):

Well, slope intercept form is y=mx+b

OpenStudy (anonymous):

What is the number you are trying to find? F or S?

OpenStudy (anonymous):

s=78 , we have to find f

OpenStudy (anonymous):

Ah, thank you! So, we need to solve for F. f=4(s-65)+10 here, we use the distributive property to get rid of the parenthesis. f=4s-260+10 and simplifying gets us f=4s-250 wala! We are now in slope intercept form!

OpenStudy (anonymous):

From this we can see that the slope of this line is 4, and the y intercept is (0, -250)

OpenStudy (anonymous):

additionally, it is easy to plug in 78 for s and find f.

OpenStudy (anonymous):

f=4(78)-250 f=312-250 f=62

OpenStudy (anonymous):

thank u but i have another problem

OpenStudy (anonymous):

ok, go!

OpenStudy (anonymous):

f=10(s-55)+40 ... s=67 & its the same way

OpenStudy (anonymous):

Well, lets do this one together.

OpenStudy (anonymous):

can you distribute out the 10?

OpenStudy (anonymous):

if s=67 then 10*67 is 670 & 10*55 is 550 .. am i rite?

OpenStudy (anonymous):

well, you want to get slope intercept first, so we don't have to deal with s=67 yet. 10(s-55) simpilifies to 10s-550

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

f=10(s-55)+40 so we now have f=10s-550 + 40

OpenStudy (anonymous):

we want to simplify it first, then plug in s.

OpenStudy (anonymous):

ohh okay! so with that wat do we do next

OpenStudy (anonymous):

after you can't simplify it any more, you are in slope intercept form, and you can plug in s.

OpenStudy (anonymous):

Continue the simplification. We have this: f=10(s-55)+40 f=10s-550 + 40 You know what to simplify next, right?

OpenStudy (anonymous):

actually no! i dont no math at all

OpenStudy (anonymous):

*know

OpenStudy (anonymous):

Well, you need to learn, and I'm afraid just giving you the answer won't help.

OpenStudy (anonymous):

So, lets proceed.

OpenStudy (anonymous):

If I have -550 + 40, how do you simpiify that?

OpenStudy (anonymous):

will it be -510

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

so the answer is f=10s-510 ... am i rite ?

OpenStudy (anonymous):

Yes! You are getting it now! f=10(s-55)+40 f=10s-550 + 40 f=10s-510 Now, we are completely simplified. It fits the form y=mx+b, or in this case f=ms+b, so we are in slope-intercept form.

OpenStudy (anonymous):

Now what?

OpenStudy (anonymous):

one min. i need tu see wat is next

OpenStudy (anonymous):

ok now! it says that f = two equations at one time

OpenStudy (anonymous):

erm, we can still plug in 67 on that other problem.

OpenStudy (anonymous):

no! this is a whole different problem

OpenStudy (anonymous):

but that is easy stuff, give me the system of equations real quick, then ill go to bed.

OpenStudy (anonymous):

lets do the system of equations real quick